OpenStudy (anonymous):

Two cyclists start biking from a trail's start 3 hours apart. The second cyclist travels at 10 miles per hour and starts 3 hours after the first cyclist who is traveling at 6 miles per hour. How much time will pass before the second cyclist catches up with the first from the time the second cyclist started biking?

OpenStudy (anonymous):

please help me this one

OpenStudy (anonymous):

s1st=s2st s1+s2=s3 6x(3+t)=10t 18=4t t=4.5 hr

OpenStudy (anonymous):

what is s1st?

OpenStudy (radar):

Did you follow that with understanding?

OpenStudy (anonymous):

not really, please explain

OpenStudy (radar):

They will catch up when they have traveled the same distance. Agree?

OpenStudy (anonymous):

yes

OpenStudy (radar):

and distance (d) is equal to speed (s) times time (t). Is that clerar?

OpenStudy (radar):

So lets set up some equations for this scenario. The slower cyclist starts out 3 hours earlier and his speed is 1/2 of the fast cyclist.

OpenStudy (radar):

Oop not quite 1/2 but slower by 4 mph lol

OpenStudy (anonymous):

The slow cyclist, traveling at 6 mph, has an 18-mile head start over the faster 10 mph biker. The biker approaches the cyclist at 4 mph (10 - 6)mph. The time required for the biker to uvertake the cyclist is (18mi)/(4 mi/h) = 4.5 hours.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what is your equation for this ?

OpenStudy (radar):

First cyclist distance will be the product of t+3 hrs and speed is 6mph. Second cyclist distance will be t times his speed of 10 mph. Set these two distance equal: 6(t+3) = 10t 6t+18=10t 4t=18 t=4.5 hours

OpenStudy (radar):

Please review all provided answers and I believe you will see the light.

OpenStudy (anonymous):

you wonderful

OpenStudy (anonymous):

i understand now dear

OpenStudy (anonymous):

Using distance = rate * time, solving for time.

OpenStudy (radar):

Good luck with your studies.

OpenStudy (anonymous):

thank you so much